# algorithmic

### v1

@@ 3 Numerical results and phase diagram
Note that there exist a number of variations of PLS, see Wegelin (2000) for a basic overview. Here we choose to compare against PLSCanonical because it treats $X$ and $Y$ symmetrically, and has higher performance than PLS-SVD as it includes the regression step from the score estimates $\hat{v}^{z}$ onto $X,Y$ to yield $\hat{w}^{z}$ as the loadings. PLS-SVD directly uses the singular vectors of $XY^{T}$ as estimates of $\hat{w}^{z}$, which performs slightly worse, see Figure 4.
For the model with all Gaussian priors, Figure 2a, we find a continuous phase transition between a tractable (easy) regime and an impossible regime. This qualitative phenomenology is the same as in the single-view case (Rangan & Fletcher, 2012; Lesieur et al., 2017). Here the algorithmic threshold obtained from Equation 29 coincides with the Bayes-optimal or information theoretic threshold, $\theta_{\mathrm{IT}}=\theta_{\mathrm{alg}}\approx 1.07$, for $\sqrt{c_{v}}=0.75$ and otherwise unit parameters. The weak recovery threshold of the rank-1 spike in each of the views $X,Y$ in isolation ($c_{v}=0$) would be $\theta^{\mathrm{single}}_{\mathrm{IT}}=1$. Therefore, a Bayes-optimal combination of information from the two modalities yields an improvement of the threshold from $\sigma_{\xi}=1$ to $\sigma_{\xi}\approx 1.07$. This improvement grows with the correlation up to $\theta_{\mathrm{IT}}\approx 1.19$ at $c_{v}=1$.
@@
For the sparse model with Rademacher-Bernoulli prior on $w^{z}$ with sparsity $\rho_{z}=0.02$ in Figure 2b, a first-order phase transition is observed instead. Again this qualitative phenomenology matches that of the single-view case (Lesieur et al., 2017), where also a technical discussion of a small regime where the lower branch acquires non-zero overlap is given. Here the algorithmic weak recovery threshold $\theta_{\mathrm{alg}}\approx 0.61$ does not coincide with the IT threshold $\theta_{\mathrm{IT}}\approx 0.71$, instead there is an algorithmically hard phase (Zdeborová & Krzakala, 2016) for $\theta_{\mathrm{alg}}<\sigma_{\xi}<\theta_{\mathrm{IT}}$ preceding the impossible regime $\sigma_{\xi}>\theta_{\mathrm{IT}}$. Again some advantage over the single-view threshold $\theta^{\mathrm{single}}_{\mathrm{alg}}\approx 0.57$ is obtained. Note that for tracing the upper branch of the phase diagram with informed AMP, we initialize the iteration at the ground-truth signal.
Finally, in Figure 5 we plot phase diagrams illustrating the algorithmic weak recovery threshold $\theta_{\mathrm{alg}}$ in the reduced three dimensional parameter space of effective snr’s $\tilde{\lambda}_{X},\tilde{\lambda}_{Y}$ and correlation coefficient $\hat{c}_{v}$. Note that $\theta_{\mathrm{IT}}$ may vary depending on the prior and is not shown, while $\theta_{\mathrm{alg}}$ is given by (29) for any zero-mean prior. Figure 5a shows for $\tilde{\lambda}_{X}=\tilde{\lambda}_{Y}$ the interpolation between zero correlation, equivalent to two single-view models, and perfect correlation, equivalent to the stackable model (20). Figure 5b illustrates the improvement of the multi-modal threshold over the thresholds of the two isolated single-view models (dashed black lines). Due to the definition of the snr (28), this plot can for example be interpreted as independently varying the aspect ratios $\alpha_{z}$. Apart from the gain in the lower left sector where no recovery is possible in any isolated model, note that also in the lower-right and upper-left sectors some degree of recovery is always possible in both modalities when the correlation is nonzero, see also Figure S2.
@@
Figure 5: Phase diagram of the algorithmic weak recovery threshold based on (29) and (30). a Varying the correlation coefficient for symmetric snr’s $\tilde{\lambda}_{X}=\tilde{\lambda}_{Y}=\tilde{\lambda}$. b The $\tilde{\lambda}_{X},\tilde{\lambda}_{Y}$ plane for $\hat{c}_{v}^{4}=1/2$. Dashed black lines show the thresholds of the modalities in isolation. The dotted line in a and b indicates the intersection of both planes.
@@ 4 Conclusions
In order to study the basic properties of multi-modal or multi-view learning, we analysed the Bayes-optimal performance of a correlated matrix factorization problem. Inferring the rank-1 spikes of the matrices corresponds to unsupervised learning of the latent variables underlying the data structure. Allowing for differences in the prior and noise channels across the two modalities or views is shown to alter the combination strategy of the AMP iteration by changing the denoising functions (11) and the $S_{z},R_{z}$ score matrices of the data. Given the combined data, the phenomenology we have observed for the Bayes-optimal learning is qualitatively the same as that of single-view learning, i.e. we have not found additional phase transitions beyond those in the single-view case.
The comparison of the Bayes-optimal weak recovery threshold and those obtained by canonical spectral methods such as PLS and CCA reveals that the canonical methods are suboptimal. This is different from the single-view case, where the optimal algorithmic weak recovery threshold agrees with the threshold present for the canonical spectral method based on principal component analysis. This difference was not anticipated by the authors nor, to our knowledge, noted in the previous literature, and it is thus worth further investigation.
In future work, it would be interesting to consider a larger number of modalities with a graph of latent relations, as in the original work of Wold (1983) and in structural equation models (Bollen, 1989). Furthermore, natural directions to explore are a supervised version of the task and how neural network-based techniques of multi-modal learning (Baltrušaitis et al., 2018) combine information from the modalities compared to the Bayes-optimal method. Both can readily be approached by considering linear or deep linear methods. An enticing question is how to share information across modalities in an approximately optimal fashion in hierarchical, non-linear models. Clues to this may well be yielded by the ongoing study of multi-sensory integration (Stein et al., 2020) in neuroscience.
Acknowledgments
We are thankful for many helpful discussions with Vittorio Erba, Odilon Duranthon and Pierre Mergny. We also thank Sagnik Nandy for our exchange about his ongoing related work, and Ilya Nemenman for a discussion which initiated the project.
@@ Appendix A Relaxed belief propagation
First we check that the peculiarity of the double product and joint prior in (8) does not cause additional correlations between the messages $\tilde{n}^{X}_{kj\to j}(v_{j})$ and $\tilde{n}^{Y}_{kj\to j}(v_{j})$ to verify that BP is applicable to this graph. This is not the case because $w^{X}_{i}$ and $w^{Y}_{i}$ are independent, so conditioned on $(v^{X}_{j},v^{Y}_{j})$, the factors $P^{X}_{\mathrm{out}}(X_{ij}|w^{X}_{i}v^{X}_{j})$ and $P^{Y}_{\mathrm{out}}(Y_{ij}|w^{Y}_{i}v^{Y}_{j})$ in (9) are not correlated; the only dependence could be inherited from $\tilde{m}^{X}_{ij\to i}(w^{X}_{i}),\;\tilde{m}^{Y}_{ij\to i}(w^{Y}_{i})$ which appear in $m^{z}_{i\to ij}(w^{z}_{i})$ both depending on $v^{z}_{j}$; however for these the structure of the factor graph is the standard dense type as, e.g. in Lesieur et al. (2017), and the dependence is sufficiently weak given a $\frac{1}{\sqrt{n}}$ scaling of the interactions. Thus (8) does not lead to additional correlations between messages that would compromise the accuracy of the BP iteration.
@@
$e^{g_{z}(z_{ij},w^{z}_{i}v^{z}_{j})}$ $=e^{g_{z}(z_{ij},0)+S^{z}_{ij}\lambda_{z}\frac{w^{z}_{i}v^{z}_{j}}{\sqrt{n_{z}}}+\frac{1}{2}(R^{z}_{ij}-(S^{z}_{ij})^{2})\lambda_{z}^{2}\frac{(w^{z}_{i}v^{z}_{j})^{2}}{n_{z}}+\mathcal{O}\left(n^{-\frac{3}{2}}_{z}\right)},$ (A.1)
$=e^{g_{z}(z_{ij},0)}\left[1+S^{z}_{ij}\lambda_{z}\frac{w^{z}_{i}v^{z}_{j}}{\sqrt{n_{z}}}+\frac{1}{2}R^{z}_{ij}\lambda_{z}^{2}\frac{(w^{z}_{i}v^{z}_{j})^{2}}{n_{z}}+\mathcal{O}\left(n^{-\frac{3}{2}}_{z}\right)\right].$ (A.2)
@@
Now to obtain rBP we use that the BP equations close on the Gaussian statistics of the messages, leading to an iteration on the means and variances of the beliefs. Plugging (A.2) into (7) (and analogously (9)) we get at the example of $\tilde{m}^{X}_{ij\to i}$
$\tilde{m}^{X}_{ij\to i}(w_{i})$ $=\frac{e^{g(X_{ij},0)}}{\mathcal{Z}^{X,m}_{ij\to i}}\int\mathrm{d}v^{X}_{j}\mathrm{d}v^{Y}_{j}\;n^{X}_{j\to ij}(v^{X}_{j},v^{Y}_{j})\left[1+S^{X}_{ij}\lambda_{X}\frac{w^{X}_{i}v^{X}_{j}}{\sqrt{n_{X}}}+\frac{1}{2}R^{X}_{ij}\lambda_{X}^{2}\frac{(w^{X}_{i}v^{X}_{j})^{2}}{n_{X}}+\mathcal{O}\left(n_{X}^{-\frac{3}{2}}\right)\right],$ (A.3)
@@
$\hat{v}^{X}_{j\to ij}$ $=\int\mathrm{d}v^{X}_{j}\mathrm{d}v^{Y}_{j}\;n^{X}_{j\to ij}(v^{X}_{j},v^{Y}_{j})v_{j}$ (A.4)
$\hat{\sigma}^{X}_{v,j\to ij}$ $=\int\mathrm{d}v^{X}_{j}\mathrm{d}v^{Y}_{j}\;n^{X}_{j\to ij}(v^{X}_{j},v^{Y}_{j})v_{j}^{2}-(\hat{v}^{X}_{j\to ij})^{2},$ (A.5)
@@
$\tilde{m}^{X}_{ij\to i}(w^{X}_{i})=\frac{1}{\mathcal{Z}^{X,m}_{ij\to i}}\exp\left[g(X_{ij},0)+S^{X}_{ij}\lambda_{X}\frac{w^{X}_{i}\hat{v}^{X}_{j\to ij}}{\sqrt{n_{X}}}-\frac{1}{2}(S^{X}_{ij})^{2}\lambda_{X}^{2}\frac{(w^{X}_{i})^{2}(\hat{v}^{X}_{j\to ij})^{2}}{n_{X}}\right.$ (A.6)
$\left.+\frac{1}{2}R^{X}_{ij}\lambda_{X}^{2}\frac{(w^{X}_{i})^{2}((\hat{v}^{X}_{j\to ij})^{2}+\hat{\sigma}^{X}_{v,j\to ij})}{n_{X}}+\mathcal{O}\left(n_{X}^{-\frac{3}{2}}\right)\right],$ (A.7)
@@
$m^{z}_{i\to ij}(w^{z}_{i})$ $=\frac{P_{w^{z}}(w^{z}_{i})}{\mathcal{Z}^{z,m}_{i\to ij}}\exp\left(J^{w}_{z,i\to ij}w^{z}_{i}-\frac{1}{2}K^{w}_{z,i\to ij}(w^{z}_{i})^{2}\right)$ (A.8)
$n^{z}_{j\to ij}(v^{X}_{j},v^{Y}_{j})$ $=\frac{P_{v}(v^{X}_{j},v^{Y}_{j})}{\mathcal{Z}^{z,n}_{j\to ij}}\exp\left(J^{v}_{z,j\to ij}v^{z}_{j}+J^{v}_{\bar{z},j}v^{\bar{z}}_{j}-\frac{1}{2}K^{v}_{z,j\to ij}(v^{z}_{j})^{2}-\frac{1}{2}K^{v}_{\bar{z},j})(v^{\bar{z}}_{j})^{2}\right)$ (A.9)
@@
$J^{v,t}_{z,j\to ij}$ $=\frac{\lambda_{z}}{\sqrt{n_{z}}}\sum_{k\neq i}^{n_{z}}S^{z}_{kj}\hat{w}^{z,t-1}_{k\to kj}$ (A.10)
$K^{v,t}_{z,j\to ij}$ $=\frac{\lambda_{z}^{2}}{n_{z}}\sum_{k\neq i}^{n_{z}}\left[(S^{z}_{kj}\hat{w}^{z,t-1}_{k\to kj})^{2}-R^{z}_{kj}((\hat{w}^{z,t-1}_{k\to kj})^{2}+\hat{\sigma}^{z,t-1}_{w,k\to kj})\right],$ (A.11)
$J^{w,t}_{z,i\to ij}$ $=\frac{\lambda_{z}}{\sqrt{n_{z}}}\sum_{k\neq j}^{d}S^{z}_{ik}\hat{v}^{z,t}_{k\to ik}$ (A.12)
$K^{w,t}_{z,i\to ij}$ $=\frac{\lambda_{z}^{2}}{n_{z}}\sum_{k\neq j}^{d}\left[(S^{z}_{ik}\hat{v}^{z,t}_{k\to ik})^{2}-R^{z}_{ik}((\hat{v}^{z,t}_{k\to ik})^{2}+\hat{\sigma}^{z,t}_{v,k\to ik})\right]$ (A.13)
where we added also explicit time indices for updating first $n^{z}_{v\to ij}(v^{z}_{j})$ then $m^{z}_{i\to ij}(w^{z}_{i})$, and the notation $J^{v}_{\bar{z},j},\;K^{v}_{\bar{z},j}$ in Equation A.9 signifies that the term $k=i$ is not excluded from the summation, which eliminates the dependence on the target node $i$. Lastly, $\hat{w}^{z,t}_{k\to kj}$ and $\hat{\sigma}^{z,t}_{w,k\to kj}$ are defined as mean and variance analogous to (A.4,A.5), but of the messages $m^{z}_{i\to ij}(w^{z}_{i})$. With the above source terms and the definition of the denoising function (11) as the first derivative of the cumulant generating function, the rBP equations with sequential update order (first $v$, then $w$) are thus
$\hat{v}^{z,t}_{j\to ij}$ $=f^{v^{z}}_{\mathrm{in}}(K^{v,t}_{X,j\to ij},\;K^{v,t}_{Y,j\to ij},\;J^{v,t}_{X,j\to ij},\;J^{v,t}_{Y,j\to ij})$ (A.14)
$\hat{\sigma}^{z,t}_{v,j\to ij}$ $=\frac{\partial f^{v^{z}}_{\mathrm{in}}}{\partial J_{z}}(K^{v,t}_{X,j\to ij},\;K^{v,t}_{Y,j\to ij},\;J^{v,t}_{X,j\to ij},\;J^{v,t}_{Y,j\to ij}).$ (A.15)
$\hat{w}^{z,t}_{i\to ij}$ $=f^{w}_{\mathrm{in}}(K^{w,t}_{z,i\to ij},J^{w,t}_{z,i\to ij})$ (A.16)
$\hat{\sigma}^{z,t}_{w,i\to ij}$ $=\frac{\partial f^{w}_{\mathrm{in}}}{\partial J_{z}}(K^{w,t}_{z,i\to ij},J^{w,t}_{z,i\to ij})$ (A.17)
While taking it into account in the following derivations, for ease of notation we have omitted in Equations A.14 and A.15 above the fact that, as explicit in Equation A.9, for $\hat{v}^{z,t+1}_{j\to ij}$ the respective $\bar{z}$ source terms do not exclude the index $i$ in the sum, $K^{v,t}_{\bar{z},j\to ij}\to K^{v,t}_{\bar{z},j}$ and $J^{v,t}_{\bar{z},j\to ij}\to J^{v,t}_{\bar{z},j}$.

### v2

@@ 4 Numerical results and comparison to PLS and CCA
Note that there exist a number of variations of PLS, see [6] for a basic overview. Here we choose to compare against PLSCanonical because it treats $X$ and $Y$ symmetrically, and has higher performance than PLS-SVD as it includes the regression step from the score estimates $\hat{v}^{z}$ onto $X,Y$ to yield $\hat{w}^{z}$ as the loadings. PLS-SVD directly uses the singular vectors of $XY^{T}$ as estimates of $\hat{w}^{z}$, which performs slightly worse, see Figure 5.
For the model with all Gaussian priors, Figure 3a, we find a continuous phase transition between a tractable (easy) regime and an impossible regime. This qualitative phenomenology is the same as in the single-view case [16, 17]. Here the algorithmic threshold obtained from Equation 29 coincides with the Bayes-optimal or information theoretic threshold, $\theta_{\mathrm{IT}}=\theta_{\mathrm{alg}}\approx 1.07$, for $\sqrt{c_{v}}=0.75$ and otherwise unit parameters. The recovery threshold of the rank-1 spike in each of the views $X,Y$ in isolation ($c_{v}=0$) would be $\theta^{\mathrm{single}}_{\mathrm{IT}}=1$. Therefore, a Bayes-optimal combination of information from the two modalities yields an improvement of the threshold from $\sigma_{\xi}=1$ to $\sigma_{\xi}\approx 1.07$. This improvement grows with the correlation up to $\theta_{\mathrm{IT}}\approx 1.19$ at $c_{v}=1$.
@@
For the sparse model with Rademacher-Bernoulli prior on $w^{z}$ with sparsity $\rho_{z}=0.02$ in Figure 3b, a first-order phase transition is observed instead. Again this qualitative phenomenology matches that of the single-view case [17], where also a technical discussion of a small regime where the lower branch acquires non-zero overlap is given. Here the algorithmic recovery threshold $\theta_{\mathrm{alg}}\approx 0.61$ does not coincide with the IT threshold $\theta_{\mathrm{IT}}\approx 0.71$, instead there is an algorithmically hard phase [13] for $\theta_{\mathrm{alg}}<\sigma_{\xi}<\theta_{\mathrm{IT}}$ preceding the impossible regime $\sigma_{\xi}>\theta_{\mathrm{IT}}$. Again some advantage over the single-view threshold $\theta^{\mathrm{single}}_{\mathrm{alg}}\approx 0.57$ is obtained. Note that for tracing the upper branch of the phase diagram with informed AMP, we initialize the iteration at the ground-truth signal.
@@ 5 Conclusions
In order to study the basic properties of multi-modal or multi-view learning, we analyzed the Bayes-optimal performance of a correlated matrix factorization problem. Inferring the rank-1 spikes of the matrices corresponds to unsupervised learning of the latent variables underlying the data structure. Allowing for differences in the prior and noise channels across the two modalities or views is shown to alter the combination strategy of the AMP iteration by changing the denoising functions (11) and the $S_{z},R_{z}$ score matrices of the data. Given the combined data, the phenomenology we have observed for the Bayes-optimal learning is qualitatively the same as that of single-view learning, i.e. we have not found additional phase transitions beyond those in the single-view case.
The comparison of the Bayes-optimal algorithmic recovery threshold and those obtained by canonical spectral methods such as PLS and CCA reveals that the canonical methods are suboptimal. This is different from the single-view case, where the optimal algorithmic recovery threshold agrees with the threshold present for the canonical spectral method based on principal component analysis. A similar necessity to reweight the entries of the naive spectral method also arises when instead of latent variables correlated across views, the noise structure of samples is correlated or heteroscedastic [51, 52, 53, 54].
In future work, it would be interesting to consider a larger number of modalities with a graph of latent relations, as in the original work of Wold [5] and in structural equation models [55]. This could be approached by starting from the matrix-tensor product model framework [42, 43]. We also note that in contrast to the recovery threshold of CCA [22], computing the threshold of PLS in random matrix theory is to our knowledge still open and could be approached using the same technique. Furthermore, natural directions to explore are a supervised version of the task and how neural network-based techniques of multi-modal learning [2] combine information from the modalities compared to the Bayes-optimal method. Both can readily be approached by considering linear or deep linear methods. An enticing question is how to share information across modalities in an approximately optimal fashion in hierarchical, non-linear models. Clues to this may well be yielded by the ongoing study of multi-sensory integration [56] in neuroscience.
We are thankful for many helpful discussions with Vittorio Erba, Odilon Duranthon and Pierre Mergny. We also thank Sagnik Nandy for our exchange about his ongoing related work, the anonymous referee for suggesting the link to the MTP model, and Ilya Nemenman for a discussion which initiated the project.
@@ Appendix A Relaxed belief propagation
First we check that the peculiarity of the double product and joint prior in (8) does not cause additional correlations between the messages $\tilde{n}^{X}_{kj\to j}(v_{j})$ and $\tilde{n}^{Y}_{kj\to j}(v_{j})$ to verify that BP is applicable to this graph. This is not the case because $w^{X}_{i}$ and $w^{Y}_{i}$ are independent, so conditioned on $(v^{X}_{j},v^{Y}_{j})$, the factors $P^{X}_{\mathrm{out}}(X_{ij}|w^{X}_{i}v^{X}_{j})$ and $P^{Y}_{\mathrm{out}}(Y_{ij}|w^{Y}_{i}v^{Y}_{j})$ in (9) are not correlated; the only dependence could be inherited from $\tilde{m}^{X}_{ij\to i}(w^{X}_{i}),\;\tilde{m}^{Y}_{ij\to i}(w^{Y}_{i})$ which appear in $m^{z}_{i\to ij}(w^{z}_{i})$ both depending on $v^{z}_{j}$; however for these the structure of the factor graph is the standard dense type as, e.g. in [17], and the dependence is sufficiently weak given a $\frac{1}{\sqrt{n}}$ scaling of the interactions. Thus (8) does not lead to additional correlations between messages that would compromise the accuracy of the BP iteration.
@@
$e^{g_{z}(z_{ij},w^{z}_{i}v^{z}_{j})}$ $=e^{g_{z}(z_{ij},0)+S^{z}_{ij}\lambda_{z}\frac{w^{z}_{i}v^{z}_{j}}{\sqrt{n_{z}}}+\frac{1}{2}(R^{z}_{ij}-(S^{z}_{ij})^{2})\lambda_{z}^{2}\frac{(w^{z}_{i}v^{z}_{j})^{2}}{n_{z}}+\mathcal{O}\left(n^{-\frac{3}{2}}_{z}\right)},$ (33)
$=e^{g_{z}(z_{ij},0)}\left[1+S^{z}_{ij}\lambda_{z}\frac{w^{z}_{i}v^{z}_{j}}{\sqrt{n_{z}}}+\frac{1}{2}R^{z}_{ij}\lambda_{z}^{2}\frac{(w^{z}_{i}v^{z}_{j})^{2}}{n_{z}}+\mathcal{O}\left(n^{-\frac{3}{2}}_{z}\right)\right].$ (34)
@@
Now to obtain rBP we use that the BP equations close on the Gaussian statistics of the messages, leading to an iteration on the means and variances of the beliefs. Plugging (34) into (7) (and analogously (9)) we get at the example of $\tilde{m}^{X}_{ij\to i}$
$\tilde{m}^{X}_{ij\to i}(w_{i})$ $=\frac{e^{g(X_{ij},0)}}{\mathcal{Z}^{X,m}_{ij\to i}}\int\mathrm{d}v^{X}_{j}\mathrm{d}v^{Y}_{j}\;n^{X}_{j\to ij}(v^{X}_{j},v^{Y}_{j})$ (35)
$\qquad\qquad\times\left[1+S^{X}_{ij}\lambda_{X}\frac{w^{X}_{i}v^{X}_{j}}{\sqrt{n_{X}}}+\frac{1}{2}R^{X}_{ij}\lambda_{X}^{2}\frac{(w^{X}_{i}v^{X}_{j})^{2}}{n_{X}}+\mathcal{O}\left(n_{X}^{-\frac{3}{2}}\right)\right],$
@@
$\hat{v}^{X}_{j\to ij}$ $=\int\mathrm{d}v^{X}_{j}\mathrm{d}v^{Y}_{j}\;n^{X}_{j\to ij}(v^{X}_{j},v^{Y}_{j})v_{j}$ (36)
$\hat{\sigma}^{X}_{v,j\to ij}$ $=\int\mathrm{d}v^{X}_{j}\mathrm{d}v^{Y}_{j}\;n^{X}_{j\to ij}(v^{X}_{j},v^{Y}_{j})v_{j}^{2}-(\hat{v}^{X}_{j\to ij})^{2},$ (37)
@@
$\tilde{m}^{X}_{ij\to i}(w^{X}_{i})=\frac{1}{\mathcal{Z}^{X,m}_{ij\to i}}\exp\left[g(X_{ij},0)+S^{X}_{ij}\lambda_{X}\frac{w^{X}_{i}\hat{v}^{X}_{j\to ij}}{\sqrt{n_{X}}}-\frac{1}{2}(S^{X}_{ij})^{2}\lambda_{X}^{2}\frac{(w^{X}_{i})^{2}(\hat{v}^{X}_{j\to ij})^{2}}{n_{X}}\right$. (38)
$\left.+\frac{1}{2}R^{X}_{ij}\lambda_{X}^{2}\frac{(w^{X}_{i})^{2}((\hat{v}^{X}_{j\to ij})^{2}+\hat{\sigma}^{X}_{v,j\to ij})}{n_{X}}+\mathcal{O}\left(n_{X}^{-\frac{3}{2}}\right)\right],$
@@
$m^{z}_{i\to ij}(w^{z}_{i})$ $=\frac{P_{w^{z}}(w^{z}_{i})}{\mathcal{Z}^{z,m}_{i\to ij}}\exp\left(J^{w}_{z,i\to ij}w^{z}_{i}-\frac{1}{2}K^{w}_{z,i\to ij}(w^{z}_{i})^{2}\right)$ (39)
$n^{z}_{j\to ij}(v^{X}_{j},v^{Y}_{j})$ $=\frac{P_{v}(v^{X}_{j},v^{Y}_{j})}{\mathcal{Z}^{z,n}_{j\to ij}}\exp\left(J^{v}_{z,j\to ij}v^{z}_{j}+J^{v}_{\bar{z},j}v^{\bar{z}}_{j}-\frac{1}{2}K^{v}_{z,j\to ij}(v^{z}_{j})^{2}-\frac{1}{2}K^{v}_{\bar{z},j})(v^{\bar{z}}_{j})^{2}\right)$ (40)
@@
$J^{v,t}_{z,j\to ij}$ $=\frac{\lambda_{z}}{\sqrt{n_{z}}}\sum_{k\neq i}^{n_{z}}S^{z}_{kj}\hat{w}^{z,t-1}_{k\to kj}$ (41)
$K^{v,t}_{z,j\to ij}$ $=\frac{\lambda_{z}^{2}}{n_{z}}\sum_{k\neq i}^{n_{z}}\left[(S^{z}_{kj}\hat{w}^{z,t-1}_{k\to kj})^{2}-R^{z}_{kj}((\hat{w}^{z,t-1}_{k\to kj})^{2}+\hat{\sigma}^{z,t-1}_{w,k\to kj})\right],$ (42)
$J^{w,t}_{z,i\to ij}$ $=\frac{\lambda_{z}}{\sqrt{n_{z}}}\sum_{k\neq j}^{d}S^{z}_{ik}\hat{v}^{z,t}_{k\to ik}$ (43)
$K^{w,t}_{z,i\to ij}$ $=\frac{\lambda_{z}^{2}}{n_{z}}\sum_{k\neq j}^{d}\left[(S^{z}_{ik}\hat{v}^{z,t}_{k\to ik})^{2}-R^{z}_{ik}((\hat{v}^{z,t}_{k\to ik})^{2}+\hat{\sigma}^{z,t}_{v,k\to ik})\right]$ (44)
where we added also explicit time indices for updating first $n^{z}_{v\to ij}(v^{z}_{j})$ then $m^{z}_{i\to ij}(w^{z}_{i})$, and the notation $J^{v}_{\bar{z},j},\;K^{v}_{\bar{z},j}$ in Equation 40 signifies that the term $k=i$ is not excluded from the summation, which eliminates the dependence on the target node $i$. Lastly, $\hat{w}^{z,t}_{k\to kj}$ and $\hat{\sigma}^{z,t}_{w,k\to kj}$ are defined as mean and variance analogous to (36,37), but of the messages $m^{z}_{i\to ij}(w^{z}_{i})$. With the above source terms and the definition of the denoising function (11) as the first derivative of the cumulant generating function, the rBP equations with sequential update order (first $v$, then $w$) are thus
$\hat{v}^{z,t}_{j\to ij}$ $=f^{v^{z}}_{\mathrm{in}}(K^{v,t}_{X,j\to ij},\;K^{v,t}_{Y,j\to ij},\;J^{v,t}_{X,j\to ij},\;J^{v,t}_{Y,j\to ij})$ (45)
$\hat{\sigma}^{z,t}_{v,j\to ij}$ $=\frac{\partial f^{v^{z}}_{\mathrm{in}}}{\partial J_{z}}(K^{v,t}_{X,j\to ij},\;K^{v,t}_{Y,j\to ij},\;J^{v,t}_{X,j\to ij},\;J^{v,t}_{Y,j\to ij}).$ (46)
$\hat{w}^{z,t}_{i\to ij}$ $=f^{w}_{\mathrm{in}}(K^{w,t}_{z,i\to ij},J^{w,t}_{z,i\to ij})$ (47)
$\hat{\sigma}^{z,t}_{w,i\to ij}$ $=\frac{\partial f^{w}_{\mathrm{in}}}{\partial J_{z}}(K^{w,t}_{z,i\to ij},J^{w,t}_{z,i\to ij})$ (48)
While taking it into account in the following derivations, for ease of notation we have omitted in Equations 45 and 46 above the fact that, as explicit in Equation 40, for $\hat{v}^{z,t+1}_{j\to ij}$ the respective $\bar{z}$ source terms do not exclude the index $i$ in the sum, $K^{v,t}_{\bar{z},j\to ij}\to K^{v,t}_{\bar{z},j}$ and $J^{v,t}_{\bar{z},j\to ij}\to J^{v,t}_{\bar{z},j}$.